# Leveraging Audio and Text Modalities in Mental Health: A Study of LLMs Performance

### v1

@@ 5.5 Multi-Label Classification Results
We also calculated the Balanced Accuracy (BA) and F1 score when treating the problem as a multiclass classification task with four classes: Depression, PTSD, both, or None. Additionally, for the multi-label classification task, we calculated BA and F1 scores based on partial correctness. For example, if the true labels were Depression and PTSD, but the model predicted only PTSD, it was credited as 50% correct since it partially matched the ground truth. This approach allows us to evaluate performance across different classification settings.
@@
Table 11: Multi label model performance on both Text and Audio modalities and their combination. Details about the specific prompts used can be found in Figure 7. The underlined bold values are the best score for the specific prompt.
@@
Audio(A) Gemini 1.5 Pro 71.9% 0.72 67.3% 0.70 67.4% 0.73 67.8% 0.74 42.5% 0.58 39.2% 0.57 70% 0.67 68% 0.60
@@
A + T Gemini 1.5 Pro 72.3% 0.71 74% 0.74 67.1% 0.72 69.8% 0.72 43% 0.56 45.9% 0.52 70% 0.72 72% 0.68
@@
In the depression detection task summarized in Table 11, Gemini 1.5 Flash achieved the best results in the text modality with a BA of 74.3% and an F1 score of 0.70 on Prompt 2. For PTSD detection, Llama 3 70Bled with a BA of 74% and an F1 score of 0.75 on Prompt 1. In the multiclass and multi-label tasks, Gemini 1.5 Flash excelled again, achieving a BA of 45.4% in the multiclass task and an F1 score of 0.81 in the multi-label task.
Gemini 1.5 Flash emerged as the most consistent model when applying the text modality, consistently achieving high scores across multiple tasks and prompts, validating its reliability in this setup.
When both modalities are combined, Gemini 1.5 Flash again shows enhanced performance, achieving a BA of 72.4% on Prompt 1 and an F1 score of 0.79 on the multi-label task under this configuration. This multimodal approach, which leverages both textual and vocal data, appears to provide a more comprehensive analysis, potentially increasing the accuracy and reliability of mental health assessments.
@@ 5.6 Disagreement
The co-occurrence matrix in Figures 9(a) and 10(a) panel visualizes the instances of correct and incorrect predictions by each modality for binary depression classification using the Gemini 1.5 Flash model on Prompt 3 and 2 respectively. The colors represent different outcomes of these predictions: Red colored cells indicate instances where both modalities incorrectly predicted the sample. Green colored cells highlight cases where both modalities correctly identified the sample. Blue colored cells denote instances where one modality outperformed the other by correctly predicting a sample that the other modality missed.
@@
Figure 9: (a) Co-occurrence matrix illustrating where the Audio and Text modalities and their combination independently predict binary depression outcomes correctly or incorrectly under zero-shot inference conditions with the Gemini 1.5 Flash model using Prompt 3. (b) Visualization of the combined (Audio + Text) modality’s predictions relative to the individual Audio and Text modalities, highlighting scenarios of agreement, disagreement, and how the combined modality addresses these differences under the same conditions.
For figure 9(a), the Modal Superiority Score (MSS) was applied to evaluate the performance differences among the modalities. The MSS metric, as detailed in Equation 4.8, measures the effectiveness of one modality over another when there is disagreement. The findings indicated that, compared to the separate modalities, the Audio+Text combined modality performed better. The advantages of modality integration were highlighted by MSS values, which demonstrated a 20% superiority of Audio+Text over Audio alone and a 30% superiority over Text alone. Audio displayed a slight advantage over Text, with an MSS value of -13.04%, indicating that Audio was marginally more successful at accurately predicting results than Text.
To further illustrate how the combined modality improved performance, we applied the Disagreement Resolvement Score (DRS) and the Modal Superiority Score (MSS) metrics on figure 9(b). The DRS, detailed in Equation 4.8, specifically addresses the combined modality’s ability to resolve conflicts between the two other modalities. It calculated a small positive value of 4.35%, indicating a slight net benefit in the combined modality’s resolution of disagreements. This suggests that the combined Audio+Text modality was able to correctly resolve just over half of the instances where Audio and Text disagreed.
Additionally, the MSS between the combined modality and the joint prediction agreement of the Audio and Text modalities showed a substantial positive score of 66.67%. This result underscores a significant enhancement in performance by the combined modality over the consensus of the individual modalities, reflecting a superior capability to effectively harness the strengths of both modalities.
@@
Using the Disagreement Resolvement Score (DRS), we evaluated the combined Audio+Text modality’s ability to resolve conflicts between the Audio and Text modalities, resulting in a significant negative score of approximately -44.83%. This significant negative score indicates that integrating Audio and Text modalities in this experiment negatively impacted the model’s ability to resolve disagreements. This finding underscores the complexities and potential challenges associated with modality integration, demonstrating that combining different modalities does not always enhance predictive accuracy.
@@ 5.7 Few-shot Prompts Results
Table 12 presents the few-shot (FS) performance scores for various tasks, where each model was subjected to tests using specifically designed prompts, as detailed in the Few-Shot Prompts section 4.6.4. The goal was to determine whether the few-shot experiments would enhance or reduce the model’s performance. Notably, for these evaluations, we utilized the entire dataset of 275 samples, including those samples that were used in the prompts. This approach was adopted because, frequently, the models failed to correctly predict at least one of these samples, thus making it crucial to include them in the performance assessment. The table lists the raw few-shot performance scores alongside the change from the zero-shot (ZS) evaluation, denoted in parentheses.
The PTSD binary task saw the largest positive shift, with Gemini 1.5 Flash (TEXT) improving by +11% in BA and +0.13 in F1. In the Depression binary task, GPT-4o mini also showed major gains, especially with a +6.6% BA improvement on Prompt 1.
Gemini 1.5 Flash (TEXT) was the most consistent, showing significant positive gains across nearly all tasks. Phi-3.5-MoE is the least consistent as it has the highest number of negative changes, including a substantial drop of -13.6% in BA for Depression severity on Prompt 2, marking it as the least consistent.
@@ 5.8 Comparative Evaluation
In this section, we undertake a comparative analysis of our model’s performance against other established methodologies, as detailed in table 13. A thorough literature search revealed a scarcity of studies that evaluated models across the entire E-DAIC or DAIC-WOZ datasets. To facilitate a robust comparison, we selected our best-performing model and the corresponding prompt. We evaluated this model in the development sets of both the E-DAIC, which contains 56 samples, and DAIC-WOZ, which consists of 35 samples. This approach allows us to directly compare our results with those obtained from other models tested under similar conditions. For our evaluation metric, we used the binary score F1, considering the task’s focus on binary depression classification, a prevalent method in existing research. We prioritized scores that were consistently high across both datasets to ensure a balanced evaluation, avoiding instances where a model performed exceptionally well on one dataset but poorly on another. This method provides a clear benchmark to assess the relative effectiveness of our proposed approaches.
@@
Table 14 offers a detailed comparative view of PTSD diagnosis performance across various methodologies, as shown below. Our zero-shot approach with GPT-4o mini is not only competitive but also surpasses the results reported by other significant studies. Notably, the CALLM model developed by Wu et al., which utilizes fine-tuning, shows a slight improvement over our approach with a balanced accuracy of 77% and an F1 score of 0.70. However, our zero-shot model still demonstrates a robust capability comparable to a finely tuned system like CALLM, outperforming other established benchmarks in the field. This underscores the potential of non-fine-tuned models to achieve high effectiveness, highlighting a significant advance in using LLMs for PTSD classification without extensive training on specific datasets.
@@
Table 12: FS model performance evaluation, Values are: FS score (Change from ZS score). The best score for each prompt is highlighted in bold and underline.
@@
Models Gemini Flash 1.5 (Text) GPT-4o mini (Text) Phi-3.5-MoE (Text) Gemini Flash 1.5 (Audio)
Task Promopt BA F1 BA F1 BA F1 BA F1
Depression Binary P1 76.1% (+2.1%) 0.65 (+0.02) 71.4% (+6.6%) 0.60 (+0.1) 71% (-3.5%) 0.60 (-0.04) 76% (+1.1%) 0.66 (0)
P2 77.4% (+4.7%) 0.67 (+0.05) 65.5% (-9%) 0.56 (-0.07) 74.7% (+1.7%) 0.64 (+0.02) 71.4% (+0.9%) 0.60 (0)
P3 77.7% (4.1%) 0.68 (+0.05) 74.2% (+9%) 0.64 (+0.14) 76.9% (+0.9%) 0.67 (0) 75.4% (+0.7%) 0.65 (+0)
PTSD Binary P1 80% ( +11%) 0.71 (+0.13) 78.2% (+1.6%) 0.70 (+0.03) 76.8% (+5.6%) 0.68 (-0.08) 75.3% (+2.4%) 0.65 (+0.03)
P2 77.5% (+7%) 0.68 (+0.7) 76% (-1%) 0.67 (-0.01) 72% (+2.8%) 0.61 (+0.05) 71.9% (+4.7%) 0.62 (+0.05)
Depression Severity P1 38.4% (+3.1%) 0.42 (+0.07) 34% (-10%) 0.32 (-0.18) 38.9% (-10%) 0.33 (-0.16) 34% (-1%) 0.42 (0)
P2 34.7% (0) 0.44 (0.14) 28.1% (-13.6%) 0.12 (-0.38) 42.9% (+10%) 0.46 (+0.04) 41.8% (+6.6%) 0.41 (+0.06)
PTSD Severity P1
LLM’s dis.:
61.2% (-1.6%)
REF dis.:
49.8% (-1.4%)
LLM’s dis.:
0.51 (-0.07)
REF dis.:
0.53 (0)
LLM’s dis.:
53.4% (-15.8%)
REF dis.:
53.5% (+2.3%)
LLM’s dis.:
0.30 (-0.29)
REF dis.:
0.50 (-0.10)
LLM’s dis.:
64.7% (+11.5%)
REF dis.:
54.4% (+2.7%)
LLM’s dis.:
0.62 (+0.23)
REF dis.:
0.49 (-0.11)
LLM’s dis.:
64.6% (+3.2%)
REF dis.:
47.8% (-3.5%)
LLM’s dis.:
0.70 (+0.14)
REF dis.:
0.54 (+0.06)
P2
LLM’s dis.:
55.4% (-1.6%)
REF dis.:
47.8% (-0.6%)
LLM’s dis.:
0.58 (+ 0.1)
REF dis.:
0.48 (+ 0.03)
LLM’s dis.:
51.1% (-17.2%)
REF dis.:
46.6% (-8%)
LLM’s dis.:
0.25 (-0.43)
REF dis.:
0.36 (-0.25)
LLM’s dis.:
59.9% (+11.3%)
REF dis.:
57.7% (+8.9%)
LLM’s dis.:
0.57 (+ 0.21)
REF dis.:
0.59 (+ 0.02)
LLM’s dis.:
65% (+4.1%)
REF dis.:
54.7% (+1.5%)
LLM’s dis.:
0.62 (+0.15)
REF dis.:
0.58 (+0.04)
Multi Label P1 71% (+3%) 0.75 (-0.06) 68% (0) 0.85 (+0.03) 73% (+8%) 0.77 (0) 69% (-1%) 0.68 (-0.03)
P2 70% (-2%) 0.75 (+0.02) 66% (-7%) 0.86 (+0.09) 72% (+7%) 0.83 (+0.14) 68% (+1%) 0.69 (-0.03)
@@
Gemini 1.5 Pro (prompt 2) (Ours) 0.56 0.69 ZS inference using the raw interview transcriptions
@@
Gemini 1.5 Pro (A+T) (prompt 3) (Ours) 0.60 0.71 ZS inference using both raw audio and transcription
@@
BERT (Danner et al.)[44] - 0.64 They trained and fine-tuned BERT for on the transcriptions. They performed ZS inference with GPT-3.5 and chatGPT-4
@@
GPT-3.5-turbo P2+SMMR (Guo et al.)[45] - 0.76 GPT-3.5-turbo and GPT-4-turbo were evaluated using ZS with a method called Stacked Multi-Model Reasoning

### v2

@@ 5.5 Multi-Label Classification Results
We also calculated the Balanced Accuracy (BA) and F1 score when treating the problem as a multiclass classification task with four classes: Depression, PTSD, both, or None. Additionally, for the multi-label classification task, we calculated BA and F1 scores based on partial correctness. For example, if the accurate labels were Depression and PTSD, but the model predicted only PTSD, it was credited as 50% correct since it partially matched the ground truth. This approach allows us to evaluate performance across different classification settings.
@@
Table 11: Multi label model performance on both Text and Audio modalities and their combination. Details about the specific prompts used can be found in Figure 7. The underlined bold values are the best score for the specific column.
@@
Deepseek_R1 70.2% 0.60 74.5% 0.70 64.9% 0.70 71.8% 0.76 43.3% 0.50 47.7% 0.63 67% 0.77 73% 0.75
Gemini Flash Lite 69.6% 0.60 70.7% 0.62 70% 0.72 68% 0.70 42.9% 0.52 43.7% 0.52 70% 0.81 69% 0.80
Gemini 2 Flash 66.7% 0.59 68.8% 0.71 66.9% 0.65 68.3% 0.70 38.1% 0.49 40.5% 0.56 67% 0.81 69% 0.70
Gemini Flash Thinking 73.4% 0.72 69.7% 0.64 64.1% 0.69 67% 0.70 39.5% 0.55 40.3% 0.52 69% 0.74 68% 0.75
Gemini 2.5 Pro 74.8% 0.71 71.8% 0.65 67% 0.72 70.7% 0.74 44.2% 0.59 43.4% 0.54 71% 0.73 71% 0.78
Llama 3.3 70B 69% 0.59 75.6% 0.69 69.8% 0.73 69.5% 0.73 41.3% 0.51 44.1% 0.59 69% 0.80 73% 0.78
Audio (A)
Gemini 1.5 Pro 71.9% 0.72 67.3% 0.70 67.4% 0.73 67.8% 0.74 42.5% 0.58 39.2% 0.57 70% 0.67 68% 0.60
@@
Gemini Flash Lite 70.7% 0.69 69.3% 0.71 71.1% 0.71 71.7% 0.75 40.4% 0.59 39% 0.61 71% 0.72 71% 0.63
Gemini 2 Flash 70.7% 0.67 71.3% 0.71 67.3% 0.71 69.7% 0.72 40.5% 0.56 45.8% 0.60 69% 0.73 71% 0.74
Gemini Flash Thinking 73.4% 0.72 71.6% 0.71 69.1% 0.75 69.2% 0.74 41.9% 0.61 44.5% 0.61 71% 0.68 70% 0.68
Gemini 2.5 Pro 72% 0.67 74.1% 0.69 69.7% 0.74 68.6% 0.72 47.9% 0.58 45.6% 0.58 71% 0.76 71% 0.76
A + T
Gemini 1.5 Pro 72.3% 0.71 74% 0.74 67.1% 0.72 69.8% 0.72 43% 0.56 45.9% 0.52 70% 0.72 72% 0.68
@@
Gemini 2 Flash 72.6% 0.69 72.1% 0.70 67.7% 0.70 64.1% 0.67 44.8% 0.56 41.7% 0.55 70% 0.77 68% 0.73
Gemini Flash Thinking 73.8% 0.68 72.6% 0.67 66.7% 0.71 69.3% 0.74 43.1% 0.55 45% 0.57 70% 0.76 71% 0.76
Gemini 2.5 Pro 70.9% 0.65 72.9% 0.66 69.9% 0.74 67.9% 0.73 43% 0.55 45.7% 0.56 70% 0.76 70% 0.77
In the depression detection task summarized in Table 11, Llama 3.3 70B achieved the best results in the text modality with a balanced accuracy (BA) of 75.6% and an F1 score of 0.69 on Prompt 2. For PTSD detection, GPT-4o mini led with a BA of 74.2% and an F1 score of 0.77 on Prompt 2. In the multiclass and multi-label tasks, Gemini 2.5 Pro achieved the highest BA of 47.9% in the multiclass task while, Phi-3.5-mini achieved an F1 score of 0.91 in the multi-label task.
Gemini 2.5 Pro emerged as the most consistent model when applying the text modality, consistently achieving high scores across multiple tasks and prompts, validating its reliability in this setup.
When both modalities are combined, Gemini 1.5 Flash again shows enhanced performance, achieving a BA of 72% on Prompt 1 and an F1 score of 0.79 on the multi-label task under this configuration. This multimodal approach, which leverages both textual and vocal data, appears to provide a more comprehensive analysis, potentially increasing the accuracy and reliability of mental health assessments.
@@
Among the evaluated models, Gemini 1.5 Flash consistently showed the strongest results in both single and multi-label tasks, particularly excelling in the text modality. GPT-4o mini also demonstrated high performance, especially in PTSD detection with Prompt 2. Deepseek_R1 presented competitive results in both multiclass and multi-label settings, especially when combining text and audio data.
Regarding the comparison between thinking and non-thinking models, the "thinking" models, such as Gemini Flash Thinking, did not consistently outperform their non-thinking counterparts. Although they occasionally achieved competitive F1 scores, like 0.72 in the audio modality, their performance varied across tasks and prompts. On the other hand, non-thinking models, especially Gemini 1.5 Pro, often provided more stable and reliable results. This suggests that the added complexity in thinking models may not consistently enhance accuracy, particularly in multi-label mental health diagnostics.
The influence of prompt design was also evident in performance variations. On average, Prompt 2 produced slightly better F1 scores across most models and tasks, particularly for PTSD and multi-label classification. However, Prompt 1 tended to yield higher balanced accuracy scores in the depression detection subtask. Prompt two may help models better identify relevant disorder cues for precision (F1), whereas Prompt 1 encourages broader recall patterns beneficial for class balance. Overall, Prompt 2 showed slightly stronger general performance, especially in complex multi-condition settings.
The analysis highlights the effectiveness of integrating text and audio modalities for mental health assessment, as combined approaches generally yielded superior results compared to single modalities. Gemini 1.5 Flash emerged as the most reliable model across diverse tasks, demonstrating consistent performance in multi-label classification scenarios. While thinking models offered some advantages, they did not consistently outperform non-thinking variants, indicating that their application should be considered based on specific diagnostic requirements. These findings underscore the importance of careful model selection and the potential benefits of multimodal approaches in complex mental health diagnostics.
@@ 5.6 Disagreement
The co-occurrence matrix in Figures 9(a) and 10(a) panels visualizes the instances of correct and incorrect predictions by each modality for binary depression classification using the Gemini 1.5 Flash model on Prompt 3 and 2, respectively. The colors represent different outcomes of these predictions: Red colored cells indicate instances where both modalities incorrectly predicted the sample. Green colored cells highlight cases where both modalities correctly identified the sample. Blue colored cells denote instances where one modality outperformed the other by correctly predicting a sample that the other modality missed.
@@
Figure 9: (a) Co-occurrence matrix illustrating where the Audio and Text modalities and their combination independently predict binary depression outcomes correctly or incorrectly under zero-shot inference conditions with the Gemini 2.5 Pro model using Prompt 1. (b) Visualization of the combined (Audio + Text) modality’s predictions relative to the individual Audio and Text modalities, highlighting scenarios of agreement, disagreement, and how the combined modality addresses these differences under the same conditions.
For figure 9(a), the Modal Superiority Score (MSS) was applied to evaluate the performance differences among the modalities. The MSS metric, as detailed in Equation 4.8, measures the effectiveness of one modality over another when there is disagreement. The findings indicated that, compared to the separate modalities, the Audio+Text combined modality performed better. The advantages of modality integration were highlighted by MSS values, which demonstrated a 52.5% superiority of Audio+Text over Audio alone and a 55.5% superiority over Text alone. Text displayed a slight advantage over Audio, with an MSS value of 11.5%, indicating that Text was marginally more successful at accurately predicting results than Audio.
To further illustrate how the combined modality improved performance, we applied the Disagreement Resolution Score (DRS) and the Modal Superiority Score (MSS) metrics on figure 9(b). The DRS, detailed in Equation 4.8, explicitly addresses the combined modality’s ability to resolve conflicts between the two other modalities. It calculated a significant positive value of 62.96%, indicating a slight net benefit in the combined modality’s resolution of disagreements. Suggests that the combined Audio+Text modality was able to correctly resolve over half of the instances where Audio and Text disagreed.
Additionally, the MSS between the combined modality and the joint prediction agreement of the Audio and Text modalities showed a substantial positive score of 43.7%. This result underscores a significant enhancement in performance by the combined modality over the consensus of the individual modalities, reflecting a superior capability to harness the strengths of both modalities effectively.
@@
Using the Disagreement Resolution Score (DRS), we evaluated the combined Audio+Text modality’s ability to resolve conflicts between the Audio and Text modalities, resulting in a significant negative score of approximately -44.83%. This significant negative score indicates that integrating Audio and Text modalities in this experiment negatively impacted the model’s ability to resolve disagreements. This finding underscores the complexities and potential challenges associated with modality integration, demonstrating that combining different modalities does not always enhance predictive accuracy.
@@ 5.7 Few-shot Prompts Results
Tables 12 present the few-shot (FS) performance scores for various tasks using specifically designed prompts, as described in Section 4.6.4. These evaluations were conducted on the full dataset of 275 samples, including those used in the prompts, due to the models’ tendency to struggle with them in zero-shot (ZS) settings. FS scores are shown alongside the change from the ZS baseline. The Gemini Flash Lite (Audio) model was excluded from these evaluations due to high variance and inconsistency in its ZS performance, which made comparative analysis unreliable.
In the binary depression task, Gemini 2.5 Pro (Audio) achieved the best result with 81.1% balanced accuracy and 0.81 F1 on Prompt 3. The most significant improvement was from Gemini Flash 1.5 (Text) on Prompt 2, with +4.7% accuracy and +0.05 F1, while Phi-3.5-MoE (Text) showed a slight decline. Audio had the highest peak, but text models were more consistent.
For binary PTSD, Gemini 2.5 Pro (Text) led with 79.9% accuracy and 0.79 F1 on Prompt 1. Gemini Flash Lite (Text) had the most substantial improvement on Prompt 2 (+11.2%, +0.07 F1). Text-based models outperformed audio, both in final scores and gains.
In depression severity, the top score came from Phi-3.5-MoE (Text) on Prompt 2 (42.9%, 0.46 F1), while Gemini Flash 1.5 (Text) improved the most on Prompt 1. GPT-4o mini had a notable drop on Prompt 2. Text remained the more stable and effective modality in this task.
For PTSD severity, Gemini 2.5 Pro (Text) on Prompt 1 reached 65% accuracy and 0.70 F1 (LLM), with strong alignment in reference distribution. It also showed the highest gain overall (+16.3%, +1.4 F1). Text models consistently outperformed audio here as well.
In the multi-label task, Gemini 2.5 Pro (Text) led with 74% accuracy and 0.81 F1 on Prompt 2, while Phi-3.5-MoE (Text) recorded the highest F1 gain (+0.14). Minor performance drops were seen in Gemini Flash Lite (Text). Text dominated this task, with multiple models surpassing 0.80 F1.
In summary, Gemini 2.5 Pro—particularly its text and audio variants—delivered the most consistent and high-performing results across tasks. Prompt 1 was most effective for PTSD binary and severity, while Prompt 3 yielded the best results for depression binary, and Prompt 2 was strongest in multi-label and severity tasks. Textual inputs generally offered broader and more stable improvements across functions, while audio demonstrated isolated peak performance, especially in depression classification. These findings demonstrate the effectiveness of tailored few-shot prompts and reinforce the potential of prompt-conditioned LLMs for scalable mental health assessment.
@@ 5.8 Comparative Evaluation
In this section, we undertake a comparative analysis of our model’s performance against other established methodologies, as detailed in tables 13. A thorough literature search revealed a scarcity of studies that evaluated models across the entire E-DAIC or DAIC-WOZ datasets. To facilitate a robust comparison, we selected our best-performing model and the corresponding prompt. We evaluated this model in the development sets of both the E-DAIC, which contains 56 samples, and DAIC-WOZ, which consists of 35 samples. This approach allows us to directly compare our results with those obtained from other models tested under similar conditions. For our evaluation metric, we used the binary score F1, considering the task’s focus on binary depression classification, a prevalent method in existing research. We prioritized scores that were consistently high across both datasets to ensure a balanced evaluation, avoiding instances where a model performed exceptionally well on one dataset but poorly on another. This method provides a clear benchmark to assess the relative effectiveness of our proposed approaches.
@@
Table 14 offers a detailed comparative view of PTSD diagnosis performance across various methodologies, as shown below. Our zero-shot approach with GPT-4o mini is not only competitive but also surpasses the results reported by other significant studies. Notably, the CALLM model developed by Wu et al., which utilizes fine-tuning, shows a slight improvement over our approach with a balanced accuracy of 77% and an F1 score of 0.70. However, our zero-shot model still demonstrates a robust capability comparable to a finely tuned system like CALLM, outperforming other established benchmarks in the field. Underscores the potential of non-fine-tuned models to achieve high effectiveness, highlighting a significant advance in using LLMs for PTSD classification without extensive training on specific datasets.
@@
Table 12: FS model performance evaluation. Values are: FS score (change from ZS score). The best score for each prompt is highlighted in bold and underlined.
@@
Task & Prompt Gemini Flash 1.5 (Text) GPT-4o mini (Text) Phi-3.5-MoE (Text) Gemini Flash 1.5 (Audio) Gemini 2.5 Pro (Text) Gemini Flash Lite (Text) Gemini 2.5 Pro (Audio)
Depression Binary P1 76.1% (+2.1%) 0.65 (+0.02) 71.4% (+6.6%) 0.60 (+0.10) 71.0% (-3.5%) 0.60 (-0.04) 76.0% (+1.1%) 0.66 (0) 78.47% (+3.6%) 0.68 (0) 76.3% (-0.6%) 0.74 (+0.01) 75.1% (-0.04%) 0.73 (+0.03)
P2 77.4% (+4.7%) 0.67 (+0.05) 65.5% (-9%) 0.56 (-0.07) 74.7% (+1.7%) 0.64 (+0.02) 71.4% (+0.9%) 0.60 (0) 74.1% (+1.1%) 0.66 (+0.01) 71.5% (+2.4%) 0.63 (+0.05) 75.9% (+0.06%) 0.65 (+0.05)
P3 77.7% (+4.1%) 0.68 (+0.05) 74.2% (+9%) 0.64 (+0.14) 76.9% (+0.9%) 0.67 (0) 75.4% (+0.7%) 0.65 (0) 79.0% (+1.75%) 0.78 (+0.01) 77.7% (+0.4%) 0.77 (+0.02) 81.1% (+2.2%) 0.81 (+0.05)
PTSD Binary P1 80.0% (+11%) 0.71 (+0.13) 78.2% (+1.6%) 0.70 (+0.03) 76.8% (+5.6%) 0.68 (-0.08) 75.3% (+2.4%) 0.65 (+0.03) 79.9% (+10.5%) 0.79 (+0.04) 67.7% (+2%) 0.56 (-0.09) 74.6% (+5.9%) 0.61 (-0.08)
P2 77.5% (+7%) 0.68 (+0.70) 76.0% (-1%) 0.67 (-0.01) 72.0% (+2.8%) 0.61 (+0.05) 71.9% (+4.7%) 0.62 (+0.05) 76.9% (+4%) 0.78 (+0.02) 76.7% (+11.2%) 0.73 (+0.07) 75.2% (+5%) 0.77 (+0.01)
Depression Severity P1 38.4% (+3.1%) 0.42 (+0.07) 34.0% (-10%) 0.32 (-0.18) 38.9% (-10%) 0.33 (-0.16) 34.0% (-1%) 0.42 (0) 44.2% (+5.2%) 0.52 (+0.06) 43.6% (+5.4%) 0.43 (+0.03) 47.3% (+6.1%) 0.53 (+0.01)
P2 34.7% (0%) 0.44 (+0.14) 28.1% (-13.6%) 0.12 (-0.38) 42.9% (+10%) 0.46 (+0.04) 41.8% (+6.6%) 0.41 (+0.06) 48.1% (+2.9%) 0.54 (-0.02) 37.5% (+3.2%) 0.40 (+0.11) 41.6% (-2%) 0.51 (+0.05)
PTSD Severity P1 LLM: 61.2% (-1.6%) LLM: 0.51 (-0.07) LLM: 53.4% (-15.8%) LLM: 0.30 (-0.29) LLM: 64.7% (+11.5%) LLM: 0.62 (+0.23) LLM: 64.6% (+3.2%) LLM: 0.70 (+0.14) LLM: 65% (+16.3%) LLM: 0.70 (+1.4) LLM: 57.3% (+4.1) LLM: 0.59 (-0.01) LLM: 62.2% (+8.8%) LLM: 0.70 (+0.10)
REF: 49.8% (-1.4%) REF: 0.53 (0) REF: 53.5% (+2.3%) REF: 0.50 (-0.10) REF: 54.4% (+2.7%) REF: 0.49 (-0.11) REF: 47.8% (-3.5%) REF: 0.54 (+0.06) REF: 58.4% (+13.2%) REF: 0.62 (+0.06) REF: 58.4% (+1.5) REF: 0.60 (-0.04) REF: 56.5% (+6.4%) REF: 0.62 (+0.01)
P2 LLM: 55.4% (-1.6%) LLM: 0.58 (+0.10) LLM: 51.1% (-17.2%) LLM: 0.25 (-0.43) LLM: 59.9% (+11.3%) LLM: 0.57 (+0.21) LLM: 65% (+4.1%) LLM: 0.62 (+0.15) LLM: 60% (+6.8%) LLM: 0.64 (+0.04) LLM: 55.1% (-3.1) LLM: 0.54 (-0.06) LLM: 55.1% (-3.1%) LLM: 0.63 (+0.01)
REF: 47.8% (-0.6%) REF: 0.48 (+0.03) REF: 46.6% (-8%) REF: 0.36 (-0.25) REF: 57.7% (+8.9%) REF: 0.59 (+0.02) REF: 54.7% (+1.5%) REF: 0.58 (+0.04) REF: 54% (-2.9) REF: 0.55 (-0.09) REF: 58.1% (-1.7) REF: 0.57 (+0.07) REF: 58.1% (-1.7%) REF: 0.66 (+0.03)
Multi Label P1 71.0% (+3%) 0.75 (-0.06) 68.0% (0%) 0.85 (+0.03) 73.0% (+8%) 0.77 (0) 69.0% (-1%) 0.68 (-0.03) 70.0% (-1%) 0.79 (+0.08) 72.0% (+2%) 0.75 (-0.06) 71.3% (+2.3%) 0.74 (+0.03)
P2 70.0% (-2%) 0.75 (+0.02) 66.0% (-7%) 0.86 (+0.09) 72.0% (+7%) 0.83 (+0.14) 68.0% (+1%) 0.69 (-0.03) 71.0% (0%) 0.81 (+0.03) 70.0% (+1%) 0.75 (-0.05) 69.8% (+2.4%) 0.76 (+0.02)
@@
Deepseek R1 (prompt 1) (T) (Ours) 0.65 0.69 ZS inference using the raw interview transcriptions
@@
Gemini 2.5 Pro (A+T) (prompt 3) (Ours) 0.60 0.71 ZS inference using both raw audio and transcription
@@
BERT (Danner et al.)[45] - 0.64
They trained and fine-tuned BERT for on the transcriptions.
They performed ZS inference with GPT-3.5 and chatGPT-4
@@
GPT-3.5-turbo P2+SMMR (Guo et al.)[46] - 0.76
GPT-3.5-turbo and GPT-4-turbo were evaluated using
ZS with a method called Stacked Multi-Model Reasoning